dr/dΩ+rTanΩ=cos2Ω
1
Expert's answer
2021-07-28T17:13:45-0400

Let us solve the differential equation "\frac{dr}{dΩ}+r\tanΩ=\cos2Ω.\\"

Firstly, let us divide both parts by "\cos\Omega." Then we have the differential equation


"\frac{1}{\cos\Omega}\frac{dr}{dΩ}+r\frac{1}{\cos\Omega}\tanΩ=\frac{1}{\cos\Omega}\cos2Ω,"


which is equivalent to the equation


"\frac{1}{\cos\Omega}\frac{dr}{dΩ}+r\frac{\sin\Omega}{\cos^2\Omega}=\frac{1}{\cos\Omega}(\cos^2Ω-\sin^2Ω),"


and hence "(r\frac{1}{\cos\Omega})'=\cosΩ-\frac{\sin^2Ω}{\cos\Omega}."


It follows that


"r\frac{1}{\cos\Omega}=\int(\cosΩ-\frac{\sin^2Ω}{\cos\Omega})d\Omega"


"=\sin\Omega-\int\frac{\sin^2Ω}{\cos^2\Omega}d(\sin\Omega)"


"=
\sin\Omega+\int\frac{\sin^2Ω}{\sin^2\Omega-1}d(\sin\Omega)"


"=
\sin\Omega+\int\frac{\sin^2Ω-1+1}{\sin^2\Omega-1}d(\sin\Omega)"


"=
\sin\Omega+\sin\Omega+\int\frac{1}{\sin^2\Omega-1}d(\sin\Omega)"


"=
2\sin\Omega+\frac{1}{2}\int(\frac{1}{\sin\Omega-1}-\frac{1}{\sin\Omega+1})d(\sin\Omega)"


"=
2\sin\Omega+\frac{1}{2}(\ln|\sin\Omega-1|-\ln|\sin\Omega+1|)+C"


"=
2\sin\Omega+\frac{1}{2}\ln|\frac{\sin\Omega-1}{\sin\Omega+1}|+C."


Therefore, the general solution is


"r= 2\sin\Omega\cos\Omega+\frac{1}{2}\cos\Omega\ln|\frac{\sin\Omega-1}{\sin\Omega+1}|+C\cos\Omega"

or

"r=
\sin2\Omega+\frac{1}{2}\cos\Omega\ln|\frac{\sin\Omega-1}{\sin\Omega+1}|+C\cos\Omega."